satellite73 (satellite73):

prove \(\sqrt{5-\sqrt{21}}=\frac{\sqrt{7}-\sqrt{3}}{\sqrt{2}}\)

OpenStudy (anonymous):

any help ? i have seen something similar. here is what wofram says http://www.wolframalpha.com/input/?i=sqrt%285-sqrt%2821%29%29

OpenStudy (callisto):

Aren't you asking the same question in different form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for two reasons. it might be easier if we know the answer and "simplify" doesn't mean anything in this context

OpenStudy (callisto):

Sorry, but one simple question: in reality, is it possible to do it backward?

OpenStudy (anonymous):

this looks for all the world like one of those trig questions where you use one half angle formula and another half angle formula and you get two different looking but equal answers. then it is a pain to show they are the same, but it usually amounts to multiplying repeatedly

OpenStudy (anonymous):

ok here is a "well known formula from basic algebra" \[\sqrt{a+b\sqrt{c}}=\sqrt{\frac{1}{2}(a+\sqrt{q})}-\sqrt{\frac{1}{2}(a-\sqrt{q})}\] if \(q=a^2-b^2c\) is a perfect square

OpenStudy (anonymous):

i must have missed that day

OpenStudy (anonymous):

guess i have my homework cut out for me

OpenStudy (anonymous):

typo actually it should be \[\sqrt{a\pm b\sqrt{c}}=\sqrt{\frac{1}{2}(a+\sqrt{q})}\pm\sqrt{\frac{1}{2}(a-\sqrt{q})}\]

OpenStudy (anonymous):

given this formula we are done, since \(5^2-21=4\) a perfect square, so we get the answer immediately. how we know this i am not sure

OpenStudy (callisto):

Sorry, I don't want to do so. but.. Please check this http://openstudy.com/study#/updates/4fda8ca0e4b0f2662fd103f2

OpenStudy (anonymous):

ok lets try this start with \[\sqrt{5-\sqrt{21}}\] and write the radicand as a perfect square. so \[\sqrt{5-\sqrt{21}}=x-y=\sqrt{(x-y)^2}=\sqrt{x^2+y^2-2xy}\] now we can put \(x^2+y^2=5, -2xy=\sqrt{21}, y=-\frac{\sqrt{21}}{2x}\) and so \(x^2+(\frac{\sqrt{21}}{2x})^2=5\) and so \[4x^4-20x^2+21=0\] solving gives \(x=\sqrt{\frac{7}{2}}\) i think

OpenStudy (anonymous):

i will have to look at this more the first "well known formula" comes directly from this paper, but no explanation, so not much help except in getting the answer

OpenStudy (anonymous):

the second method i tried to mimic what i read here http://gauravtiwari.org/2011/03/16/a-problem-on-ordinary-nested-radicals/ but the line "which on simplification yields..." was not so obvious to me

OpenStudy (anonymous):

ok, take a look at 4.5.2 here and the following theorem. this i think give it all

OpenStudy (anonymous):

learn something new every day!

OpenStudy (anonymous):

today for example i learned where to get a really great hamburger

OpenStudy (anonymous):

your friend must have run in to someone who was either nuts ( a nutty math teacher, imagine) or just learned this and wanted to show off. i cannot imagine this in a basic algebra class

OpenStudy (callisto):

*Disappointed* You're not learning HOW to get a great hamburger but you learnt WHERE to get it :( Thanks for all the notes and the 'well unknown formula' which I haven't heard of. Hmm.. Actually, I can solve it :|

OpenStudy (anonymous):

great. let me know how it goes. if they had had the internet when i was in school i would have 4 degrees by now

OpenStudy (kinggeorge):

This might be good enough. Look at \(\sqrt{5-\sqrt{21}}\). If you square it, you get \(5-\sqrt{21}\). We want this to be a perfect square of some binomial with square roots. This means, that we should want a square multiplied by two. Hence, multiply this by two. We get \[10-2\sqrt{21}\]If we want this to be a perfect square of some expression that looks like \(\sqrt x\pm\sqrt y\), we want to find an \(x, y\) such that \(x+y=10\) and \(xy=21\). Solving this, we get that \[10-2\sqrt{21}=(\sqrt7-\sqrt3)^2\]Now we work backward to get what we had at the start. Divide by 2, and then take the square root gives us that \[\sqrt{5-\sqrt{21}}=\frac{\sqrt7-\sqrt3}{\sqrt2}\]

OpenStudy (kinggeorge):

Typo in that first paragraph. I said "This means, that we should want a square multiplied by two." It should read "This means, that we should want a square root multiplied by two." (talking about the square root in \(5-\sqrt{21}\) of course)

OpenStudy (callisto):

@KingGeorge I solved the question using this way. But no one here looks at it except myininaya

OpenStudy (anonymous):

this looks much snappier

OpenStudy (kinggeorge):

The two methods are very slightly different. Same trick, but slightly different process.

OpenStudy (kinggeorge):

Namely, I took things out of the square root, and you left them in. So I guess they really are just the same thing.